Circular No. 7980

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2001 U 1 A new candidate for a satellite of Uranus, of mag R = 25.3–25.5, was

found by M. Holman, Harvard-Smithsonian Center for Astrophysics (CfA); J. Kavelaars, now at Dominion Astrophysical Observatory; and D. Milisavljevic, McMaster University, with the 4-m reflector at the Cerro Tololo Interamerican Observatory (CTIO) on 2001 Aug. 13. Kavelaars obtained confirmatory images with the Canada-France-Hawaii Telescope on Aug. 25. Following further measurements by B. Gladman, from images obtained by P. Nicholson and C. Dumas with the 5-m Hale reflector at Palomar on Sept. 21 and by J.-M. Petit with the European Southern Observatory 8.2-m UT3 telescope at Paranal on Nov. 15, B. G. Marsden, CfA, provided a prediction for the 2002 opposition that allowed Gladman to recover a candidate for the object on Paranal images obtained by P. Rousselot and O. Mousis on 2002 Sept. 5, after which Holman, with the help of T. Grav, confirmed the recovery on CTIO 4-m reflector images obtained on 2002 Aug. 13. Marsden established definitively an orbit with a = 0.057 AU, e = 0.21, i = 166^{o} (to the ecliptic) and that H = 12.8. Full details of the astrometry and orbit calculation, as well as a current ephemeris, are provided on MPEC 2002-S64.

S/2002 (121) 1 W. J. Merline, Southwest Research Institute (SwRI); P. M. Tamblyn,

Binary Astronomy and SwRI; C. Dumas, Jet Propulsion Laboratory (JPL); L. M. Close, University of Arizona; C. R. Chapman, SwRI; F. Menard, Observatoire de Grenoble; W. M. Owen, JPL; and D. C. Slater, SwRI; and J. Pepin, University of Colorado, report the discovery on Sept. 28.6 UT of a satellite of (121) Hermione (a C-type, Cybele-family asteroid of diameter 209 km) from J-, H-, and K′-band direct imaging with the 10-m Keck II Telescope (+ NIRC2/AO adaptive optics sysem) on Mauna Kea. The satellite was tracked until it became inaccessible to the telescope; over this 0.5-hr baseline, no relative motion was detected (a search for known background/foreground small bodies with similar position and velocity reveals no candidates closer than 1^{o} that are of similar brightness to the satellite). On Sept. 28.6334, the secondary was at 0′′.43 separation (projected separation 630 km) in p.a. 257^{o}. The K brightness difference is ∼ 6 mag, giving an estimated diameter of the satellite of ∼ 13 km.

2002 September 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green